OpenStudy (anonymous):

A dripping faucet waste a cup of water if it drips for 4 minutes the amount of water wasted varies directly with the amount of time the faucet drips what is the constant of variation

OpenStudy (danjs):

'the amount of water wasted varies directly with the amount of time the faucet drips' in other words, directly proportional , if time increases, so does water wasted \[\large water~ wasted [w]~~~\alpha~~~time~running[t]\] \[\Large W~~\alpha~~t\]

OpenStudy (danjs):

To make an equation , you need a proportion constant, k W = k * t They gave you one point ( t , W) = (4min , 1 cup), use that to figure the constant k